OpenStudy (anonymous):

find the total area of the cylinder given the radius(r) or diameter(d) and height (h) d=18ft h=3.5m

OpenStudy (anonymous):

Area is twice the area of the circular caps plus the lateral surface area. \[2*\frac{1}{4}\pi 18^2+\pi 18*\frac{35}{10}=225\pi \text{ \square meters} \]